OpenStudy (firejay5):

The sum of the measures of the angles of a polygon with n sides is given. Find n. 32. 180 35. 2,880 Show work/explain answer of how you got it for you to receive a medal for your patronage! :D

OpenStudy (anonymous):

Use the formula, the sum of the interior angles of a polygon is: 180(n - 2) where n + # of sides the polygon has So just substitute to find n.

OpenStudy (firejay5):

substitute 180 for n

OpenStudy (anonymous):

No, n represents the number of sides. Look....... Total number of degrees in the interior angles of a polygon = 180(n-2)

OpenStudy (anonymous):

So let's say...the total number of degrees is 2,880. So your equation is: 2,880 = 180(n - 2)

OpenStudy (anonymous):

Divide both sides by 180, you get 16 = n - 2 add 2 to both sides, you get n = 18.

OpenStudy (firejay5):

I got it now! :D @Easyaspi314 Thanks dude! :D

OpenStudy (solomonzelman):

Nice!

OpenStudy (anonymous):

welcome.